# Maximum Likelihood Estimation Based Complex-Valued Robust Chinese Remainder Theorem and Its Fast Algorithm

### v1

@@ IV-B Probability of Error Preserving MLE C-CRT
We now calculate the probability of the MLE C-CRT preserving errors, i.e., satisfying the necessary and sufficient conditions in Theorem 6. This is also a probability for achieving robust reconstruction when $\tau=\frac{M}{2}$ by Corollary 2. Assume that the $i$-th remainder error $\Delta\mathsf{r}_{i}$ follows a wrapped complex Gaussian distribution with a mean of 00 and a variance of $2\sigma_{i}^{2}$, and that the real and imaginary parts of $\Delta\mathsf{r}_{i}$ have equal variance $\sigma_{i}^{2}$. Since $|M\Gamma_{i}|$ is generally much larger than $\sigma_{i}^{2}$, we approximate $\Delta\mathsf{r}_{i}$ as a complex Gaussian distribution.
@@
$p((\Delta\mathsf{r}_{1},\Delta\mathsf{r}_{2},\ldots,\Delta\mathsf{r}_{L})\in\mathcal{R})$
$=$ $\textstyle\frac{1}{(2\pi)^{L}}\prod_{i=1}^{L}\frac{1}{\sigma_{i}^{2}}\underset{\mathbf{r}\in\mathcal{R}}{\idotsint}\exp{\left\{\sum_{i=1}^{L}\left(-\frac{x_{i}^{2}}{2\sigma_{i}^{2}}-\frac{y_{i}^{2}}{2\sigma_{i}^{2}}\right)\right\}}\mathrm{d}V_{\mathbf{x}}\mathrm{d}V_{\mathbf{y}}$
$=$ $\textstyle\frac{1}{(2\pi)^{L}}\prod_{i=1}^{L}\frac{1}{\sigma_{i}^{2}}\bigg{(}\underset{\mathbf{r}\in\mathcal{R}}{\idotsint}\exp\left\{-\sum_{i=1}^{L}\frac{x_{i}^{2}}{2\sigma_{i}^{2}}\right\}\mathrm{d}V_{\mathbf{x}}\bigg{)}^{2},$
@@ V-A Comparison of Fast MLE C-CRT and Two-Stage CRT Algorithms
In this subsection we compare the proposed fast MLE C-CRT and the two-stage CRT presented in [7] in terms of both performance and computational complexity, where the moduli $\mathsf{\Gamma}_{1},\mathsf{\Gamma}_{2},\ldots,\mathsf{\Gamma}_{8}$ are set as $71+4\mathrm{i},1-4\mathrm{i},3+4\mathrm{i},3-4\mathrm{i},2+7\mathrm{i},2-7\mathrm{i},3,$ 7, respectively. In each trial, the real and imaginary parts of the complex number $\mathsf{N}$ are randomly selected from the interval $[M,M(\Gamma-1))$, where $M=10$. The real and imaginary parts of the remainder errors $\Delta\mathsf{r}_{i}$ follow a wrapped complex Gaussian distribution with mean 00 and variance $\sigma_{i}^{2}$. In the simulation, we set $\sigma_{i}=u|M\mathsf{\Gamma}_{i}|$, where $u$ is a small positive constant. For convenience, we approximate $\mathsf{r}_{i}$ as a uniform distribution within $\mathcal{F}_{M\mathsf{\Gamma}_{i}}$. Similar to (45), $-20\log_{10}\sqrt{3}u$ can be used as the measurement for SNR of the remainders. For each $u$, the total number $n$ of trials is $10000$, i.e., $n=10000$. We evaluate the performance of the two methods using two metrics: the root mean square error (RMSE), and the trial fail rate (TFR) for robust reconstruction and preserving errors. The RMSE of $\mathsf{N}$ is defined as
@@
where $E\{\cdot\}$ denotes the mean. Since $\mathrm{Re}(\Delta\mathsf{r}_{i})$ are mutually independent and Gaussian distributed for $i=1,2,\ldots,L$, $\mathrm{Re}(\overline{\Delta\mathsf{r}})$ follows a Gaussian distribution with mean 00 and variance $\sum_{i=1}^{L}w_{i}^{2}\sigma_{i}^{2}$. Similarly, the distribution of $\mathrm{Im}(\overline{\Delta\mathsf{r}})$ is the same as that of $\mathrm{Re}(\overline{\Delta\mathsf{r}})$. Thus,
@@
Fig. 9 illustrates the curves of the RMSE for the two algorithms in terms of SNR, along with the theoretical RMSE of the fast MLE C-CRT. It can be observed that the RMSE of the fast MLE C-CRT is smaller than that of the two-stage CRT. At the SNR of $32$dB, the maximum error in the real and imaginary parts of the remainders is $4.3267$, the reconstruction errors for the fast MLE C-CRT are less than $1.2621$. At the SNR of $34$dB, the maximum error in the remainders is $3.1640$, the reconstruction errors for the two-stage CRT are less than $1.1697$. From Fig. 9, one can see that the fast MLE C-CRT achieves robust reconstruction more quickly than the two-stage CRT does, since it has the optimal estimation of the common remainder. When the SNR is less than $30$dB, due to the errors of some remainders exceeding $\frac{M}{2}$, this does not satisfy the conditions of Theorem 6. Hence, the fast MLE C-CRT may not have robust reconstruction and has large errors. Similarly, when the SNR is less than $34$dB, the condition for the robust reconstruction of the two-stage CRT algorithm may not be satisfied and thus has large errors. On the other hand, the theoretical curve is from (46) that is based on the assumption of complex Gaussian distributions of the remainder errors and only depends on the error distribution variances, while the true distributions of the remainder errors follow wrapped complex Gaussian distributions. It justifies that the theoretical curve is smooth and does not exhibit a large change. When the SNR is higher, the assumption holds better and the simulated curve and the theoretical curve match better. As one can see from the zoom-in part in Fig. 9, the fast MLE C-CRT, in fact, achieves the theoretical RMSE values when SNR is high, while the two-stage CRT cannot.
@@
For the TFR in terms of preserving errors, we test whether (44) holds. If it does, the trial is considered successful and otherwise, the trial fails. According to Theorem 6, the TFR of preserving errors for each trial can be expressed as $p_{\text{TFR}}^{L}=1-p((\Delta\mathsf{r}_{1},\Delta\mathsf{r}_{2},\ldots,\Delta\mathsf{r}_{L})\in\mathcal{R})$. Fig. 11 illustrates the curves of the TFR and its theoretical value for the fast MLE C-CRT with respect to the number of moduli $L$, where the noise variances $2\sigma_{i}^{2}$ are constant and equal to $2\sigma^{2}$ for $i=1,2,\ldots,L$. It can be observed that the TFR of the fast MLE C-CRT closely matches its theoretical value in both cases.
@@
We now compare the computational complexities of the two methods by counting the numbers of the real multiplications they need, i.e., the real multiplicative complexities. Given a known common remainder, the real multiplicative complexity for the fast MLE C-CRT can be directly calculated as $\mathcal{O}(L)$ by (28), (29), and (30). The two-stage CRT requires twice the number of operations of the closed-form CRT for reals, with a real multiplicative complexity of $\mathcal{O}(L)$ when the common remainder is known [11]. Therefore, the computational complexities of the two algorithms primarily arise from the estimation of the common remainder $\hat{\mathsf{r}}^{c}$. For the fast MLE C-CRT, $\hat{\mathsf{r}}^{c}$ is obtained from the objective functions in (33). Each evaluation of the real or imaginary part requires $4L$ real multiplications. According to Theorem 4, obtaining $\hat{\mathsf{r}}^{c}$ necessitates $8L^{2}$ real multiplications in total. When $L\geq 5$, the common remainder search in the two-stage CRT arises from its second stage, where the objective functions are the special case of (33) with $w_{i}=1$ for each $i$. According to the algorithm presented in [11], each evaluation of the real or imaginary part involves $5(L-l)$ real multiplications, where $2l\leq L$ is the number of complex-valued moduli. Denoting $\epsilon$ as the search step size, this algorithm requires at least $\frac{5ML}{\epsilon}$ real multiplications to estimate the common remainder. Since $\frac{M}{\epsilon}$ is generally much larger than $L$, the complexity of the two-stage CRT is higher in this case. As mentioned earlier, the two-stage CRT proposed in [7] requires a search process to estimate the common remainder in [11] but does not utilize the fast algorithm proposed in [21]. If so, the number of searches would be $2(L-l)$ and only $10(L-l)^{2}$ real multiplications are required to obtain the common remainder. Consequently, the reference remainder can be properly determined and hence the folding integers ($n_{i}$ defined in [11]) can be correctly determined. However, the estimation is not the MLE since the estimation of the common remainder can not be utilized in the reconstruction of $\mathsf{N}$. Therefore, although the two-stage CRT provides a robust estimation, it is not optimal even when utilizing the fast algorithm proposed in [21].
@@ V-B Application in Modulo ADCs
The condition for the MLE C-CRT to uniquely reconstruct $\mathsf{N}$ from the congruence system (9) is that $\mathsf{N}\in\mathcal{F}_{M\Gamma}$, which is the same as that of the two-stage CRT in the sense of congruence. Specifically, if the moduli in (9) are the real numbers $M\Gamma_{1},M\Gamma_{2},\ldots,M\Gamma_{L}$, then we have $\mathrm{Re}(\mathsf{N})\equiv\mathrm{Re}(\mathsf{r}_{i})\mod M\Gamma_{i},\ i=1,2,\ldots,L$, and $\mathrm{Im}(\mathsf{N})\equiv\mathrm{Im}(\mathsf{r}_{i})\mod M\Gamma_{i},\ i=1,2,\ldots,L$. If $\mathrm{Re}(\mathsf{N})$ and $\mathrm{Im}(\mathsf{N})$ are within $[0,M\Gamma)$, then $\mathsf{N}$ can be uniquely reconstructed by the two independent real-valued CRTs.
For an integer $M\Gamma$ while $\mathsf{\Gamma}_{i},i=1,2,\ldots,L$, are Gaussian integers, we know that $\mathcal{F}_{M\Gamma}$ is a square according to (2) as also mentioned earlier, where its sides are parallel to the real and imaginary axes with length $M\Gamma$. From Theorem 1, any complex number in $\mathcal{F}_{M\Gamma}$ can be uniquely reconstructed by C-CRT, i.e., the multi-channel SR-ADCs in Fig. 3. Thus, both of the uniquely determinable real and imaginary parts of a complex number by using C-CRT or the multi-channel SR-ADCs are within $[0,M\Gamma)$. In this case, we call $M\Gamma$ as the dynamic range of the C-CRT. Thus, within this dynamic range, the sampled values of $\mathsf{g}(t)$ can be uniquely recovered by the C-CRT and the two-stage CRT with complex-valued moduli, or the closed-form CRTs with real-valued moduli. Note that to have an integer lcm, the complex-valued moduli can be selected as pairs of conjugate Gaussian integers.
@@
$\Delta_{\max}$ Method $M\mathsf{\Gamma}_{i}$ Dynamic
range $M\Gamma$
@@ 9 C-CRT, Two-stage CRT $7+4\mathrm{i}$, $7-4\mathrm{i}$, 9 585
where $A$ is a constant, coefficients $a_{k}$ and $b_{k}$ are uniformly distributed in $[-1,1]$, $\mathrm{sinc}(t)=\frac{\sin(\pi t)}{\pi t}$. For the closed-form CRTs, we set six SR-ADCs with dynamic ranges of $5,5,6,6,7$, and $7$ when $\Delta_{\max}=7$, and $7,7,8,8,9$, and $9$ when $\Delta_{\max}=9$. For the C-CRT and the two-stage CRT, we set six SR-ADCs with dynamic ranges of $\sqrt{41}$, $\sqrt{41}$, $\sqrt{41}$, $\sqrt{41}$, $7$ and $7$ when $\Delta_{\max}=7$, and $\sqrt{65}$, $\sqrt{65}$, $\sqrt{65}$, $\sqrt{65}$, $9$ and $9$ when $\Delta_{\max}=9$, where the moduli $M\mathsf{\Gamma}_{i}$ are shown in Table I. The reconstruction error is quantified using the root relative squared error (RRSE), given by
@@
where the sampling frequency $f_{s}$ in the time domain for each channel is the Nyquist rate, i.e., $1$Hz.
Fig. 13 illustrates the RRSE curves for the three methods in terms of SNR. The fast MLE C-CRT demonstrates the best performance overall. Similar to the previous RMSE simulations, the fast MLE C-CRT achieves robustness at an SNR of $16$dB. In contrast, the two-stage CRT achieves robustness at an SNR of $18$dB. Additionally, for SNR values of $18$dB and higher, the performances of the two-stage CRT and the closed-form CRT are nearly indistinguishable.
@@
To show the high dynamic range of complex moduli, while the maximum dynamic range of all the SR-ADCs is constrained by $\Delta_{\max}$, we compute the TFR of the robust reconstruction for the three methods. The curves of the TFR for the three methods in terms of SNR are illustrated in Fig. 14, where $a_{k}=-0.9$, $b_{k}=0.98$, and $\tau=0.25$. Since some sampled values exceed the dynamic range of SR-ADCs with real-valued moduli, the closed-form CRTs exhibit an error floor. Additionally, although the TFRs of both the fast MLE C-CRT and the two-stage CRT can approach 00, the fast MLE C-CRT outperforms the two-stage CRT due to its optimal estimation of the common remainder. These results illustrate that the proposed MLE C-CRT for complex numbers performs better than the conventional CRT for real values. In addition, if C-CRT is thought of as a special 2D-CRT, it clearly shows that 2D-CRT (non-separable) performs better than two 1D-CRTs (separable).

### v3

@@ IV-B Probability of Error Preserving MLE C-CRT
We now calculate the probability of the MLE C-CRT preserving errors, i.e., satisfying the necessary and sufficient conditions in Theorem 6. This is also a probability for achieving robust reconstruction when $\tau=\frac{M}{2}$ by Corollary 2. Assume that the $i$-th remainder error $\Delta\mathsf{r}_{i}$ follows a wrapped complex Gaussian distribution with a mean of 0 and a variance of $2\sigma_{i}^{2}$, and that the real and imaginary parts of $\Delta\mathsf{r}_{i}$ have equal variance $\sigma_{i}^{2}$. Since $|M\Gamma_{i}|$ is generally much larger than $\sigma_{i}^{2}$, we approximate $\Delta\mathsf{r}_{i}$ as a complex Gaussian distribution.
@@
$\hskip-3.00003pt\begin{aligned} &p((\Delta\mathsf{r}_{1},\Delta\mathsf{r}_{2},\ldots,\Delta\mathsf{r}_{L})\in\mathcal{R})\\ =&\frac{1}{(2\pi)^{L}}\prod_{i=1}^{L}\frac{1}{\sigma_{i}^{2}}\underset{\mathbf{r}\in\mathcal{R}}{\idotsint}\exp{\left\{\sum_{i=1}^{L}\left(-\frac{x_{i}^{2}}{2\sigma_{i}^{2}}-\frac{y_{i}^{2}}{2\sigma_{i}^{2}}\right)\right\}}\mathrm{d}V_{\mathbf{x}}\mathrm{d}V_{\mathbf{y}}\\ =&\frac{1}{(2\pi)^{L}}\prod_{i=1}^{L}\frac{1}{\sigma_{i}^{2}}\bigg{(}\underset{\mathbf{r}\in\mathcal{R}}{\idotsint}\exp\left\{-\sum_{i=1}^{L}\frac{x_{i}^{2}}{2\sigma_{i}^{2}}\right\}\mathrm{d}V_{\mathbf{x}}\bigg{)}^{2},\end{aligned}$
@@ V-A Comparison of Fast MLE C-CRT and Two-Stage CRT Algorithms
In this subsection we compare the proposed fast MLE C-CRT and the two-stage CRT presented in [23] in terms of both performance and computational complexity, where the moduli $\mathsf{\Gamma}_{1},\mathsf{\Gamma}_{2},\ldots,\mathsf{\Gamma}_{8}$ are set as $1+4\mathrm{i},1-4\mathrm{i},3+4\mathrm{i},3-4\mathrm{i},2+7\mathrm{i},2-7\mathrm{i},3,7$, respectively. In each trial, the real and imaginary parts of the complex number $\mathsf{N}$ are randomly selected from the interval $M,M(\Gamma-1))$, where $M=10$. The real and imaginary parts of the remainder errors $\Delta\mathsf{r}_{i}$ follow a wrapped complex Gaussian distribution with mean 0 and variance $\sigma_{i}^{2}$. In the simulation, we set $\sigma_{i}=u|M\mathsf{\Gamma}_{i}|$, where $u$ is a small positive constant. For convenience, we approximate $\mathsf{r}_{i}$ as a uniform distribution within $\mathcal{F}_{M\mathsf{\Gamma}_{i}}$. Similar to (46), $20\log_{10}\sqrt{3}u$ can be used as the measurement for SNR of the remainders. For each $u$, the total number $n$ of trials is $10000$, i.e., $n=10000$. We evaluate the performance of the two methods using two metrics: the root mean square error (RMSE), and the trial fail rate (TFR) for robust reconstruction and preserving errors. The RMSE of $\mathsf{N}$ is defined as
@@
where $E\{\cdot\}$ denotes the mean. Since $\mathrm{Re}(\Delta\mathsf{r}_{i})$ are mutually independent and Gaussian distributed for $i=1,2,\ldots,L$, $\mathrm{Re}(\overline{\Delta\mathsf{r}})$ follows a Gaussian distribution with mean 0 and variance $\sum_{i=1}^{L}w_{i}^{2}\sigma_{i}^{2}$. Similarly, the distribution of $\mathrm{Im}(\overline{\Delta\mathsf{r}})$ is the same as that of $\mathrm{Re}(\overline{\Delta\mathsf{r}})$. Thus,
@@
Fig. 9 illustrates the curves of the RMSE for the two algorithms in terms of SNR, along with the theoretical RMSE of the fast MLE C-CRT. It can be observed that the RMSE of the fast MLE C-CRT is smaller than that of the two-stage CRT. At the SNR of 3232dB, the maximum error in the real and imaginary parts of the remainders is $4.3267$, the reconstruction errors for the fast MLE C-CRT are less than $1.2621$. At the SNR of 3434dB, the maximum error in the remainders is $3.1640$, the reconstruction errors for the two-stage CRT are less than $1.1697$. From Fig. 9, one can see that the fast MLE C-CRT achieves robust reconstruction more quickly than the two-stage CRT does, since it has the optimal estimation of the common remainder. When the SNR is less than 3030dB, due to the errors of some remainders exceeding $\frac{M}{2}$, this does not satisfy the conditions of Theorem 6. Hence, the fast MLE C-CRT may not have robust reconstruction and has large errors. Similarly, when the SNR is less than 3434dB, the condition for the robust reconstruction of the two-stage CRT algorithm may not be satisfied and thus has large errors. On the other hand, the theoretical curve is from (47) that is based on the assumption of complex Gaussian distributions of the remainder errors and only depends on the error distribution variances, while the true distributions of the remainder errors follow wrapped complex Gaussian distributions. It justifies that the theoretical curve is smooth and does not exhibit a large change. When the SNR is higher, the assumption holds better and the simulated curve and the theoretical curve match better. As one can see from the zoom-in part in Fig. 9, the fast MLE C-CRT, in fact, achieves the theoretical RMSE values when SNR is high, while the two-stage CRT cannot.
@@
For the TFR in terms of preserving errors, we test whether (45) holds. If it does, the trial is considered successful and otherwise, the trial fails. According to Theorem 6, the TFR of preserving errors for each trial can be expressed as
$p_{\text{TFR}}^{L}=1-p((\Delta\mathsf{r}_{1},\Delta\mathsf{r}_{2},\ldots,\Delta\mathsf{r}_{L})\in\mathcal{R}).$
Fig. 11 illustrates the curves of the TFR and its theoretical value for the fast MLE C-CRT with respect to the number of moduli $L$, where the noise variances $2\sigma_{i}^{2}$ are constant and equal to $2\sigma^{2}$ for $i=1,2,\ldots,L$. It can be observed that the TFR of the fast MLE C-CRT closely matches its theoretical value in both cases.
@@
We now compare the computational complexities of the two methods by counting the numbers of the real multiplications they need, i.e., the real multiplicative complexities. Given a known common remainder, the real multiplicative complexity for the fast MLE C-CRT can be directly calculated as $\mathcal{O}(L)$ by (29), (30), and (31). The two-stage CRT requires twice the number of operations of the closed-form CRT for reals, with a real multiplicative complexity of $\mathcal{O}(L)$ when the common remainder is known [7]. Therefore, the computational complexities of the two algorithms primarily arise from the estimation of the common remainder $\hat{\mathsf{r}}^{c}$. For the fast MLE C-CRT, $\hat{\mathsf{r}}^{c}$ is obtained from the objective functions in (34). Each evaluation of the real or imaginary part requires $4L$ real multiplications. According to Theorem 4, obtaining $\hat{\mathsf{r}}^{c}$ necessitates $8L^{2}$ real multiplications in total. When $L\geq 5$, the common remainder search in the two-stage CRT arises from its second stage, where the objective functions are the special case of (34) with $w_{i}=1$ for each $i$. According to the algorithm presented in [7], each evaluation of the real or imaginary part involves 5(Ll)5(L-l) real multiplications, where $2l\leq L$ is the number of complex-valued moduli. Denoting $\epsilon$ as the search step size, this algorithm requires at least $\frac{5ML}{\epsilon}$ real multiplications to estimate the common remainder. Since $\frac{M}{\epsilon}$ is generally much larger than $L$, the complexity of the two-stage CRT is higher in this case. As mentioned earlier, the two-stage CRT proposed in [23] requires a search process to estimate the common remainder in [7] but does not utilize the fast algorithm proposed in [10]. If so, the number of searches would be 2(Ll)2(L-l) and only $10(L-l)^{2}$ real multiplications are required to obtain the common remainder. Consequently, the reference remainder can be properly determined and hence the folding integers ($n_{i}$ defined in [7]) can be correctly determined. However, the estimation is not the MLE since the estimation of the common remainder can not be utilized in the reconstruction of $\mathsf{N}$. Therefore, although the two-stage CRT provides a robust estimation, it is not optimal even when utilizing the fast algorithm proposed in [10].
@@ V-B Application in Modulo ADCs
The condition for the MLE C-CRT to uniquely reconstruct $\mathsf{N}$ from the congruence system (10) is that $\mathsf{N}\in\mathcal{F}_{M\Gamma}$, which is the same as that of the two-stage CRT in the sense of congruence. Specifically, if the moduli in (10) are the real numbers $M\Gamma_{1},M\Gamma_{2},\ldots,M\Gamma_{L}$, then we have
$\mathrm{Re}(\mathsf{N})\equiv\mathrm{Re}(\mathsf{r}_{i})\mod M\Gamma_{i},\ i=1,2,\ldots,L,$
and
$\mathrm{Im}(\mathsf{N})\equiv\mathrm{Im}(\mathsf{r}_{i})\mod M\Gamma_{i},\ i=1,2,\ldots,L.$
If $\mathrm{Re}(\mathsf{N})$ and $\mathrm{Im}(\mathsf{N})$ are within $0,M\Gamma)$, then $\mathsf{N}$ can be uniquely reconstructed by the two independent real-valued CRTs.
For an integer $M\Gamma$ while $\mathsf{\Gamma}_{1},\mathsf{\Gamma}_{2},\ldots,\mathsf{\Gamma}_{L}$ are Gaussian integers, we know that $\mathcal{F}_{M\Gamma}$ is a square according to (2) as also mentioned earlier, where its sides are parallel to the real and imaginary axes with length $M\Gamma$. From Theorem 1, any complex number in $\mathcal{F}_{M\Gamma}$ can be uniquely reconstructed by C-CRT, i.e., the multi-channel SR-ADCs in Fig. 3. Thus, both of the uniquely determinable real and imaginary parts of a complex number by using C-CRT or the multi-channel SR-ADCs are within $0,M\Gamma)$. In this case, we call $M\Gamma$ as the dynamic range of the C-CRT. Thus, within this dynamic range, the sampled values of $\mathsf{g}(t)$ can be uniquely recovered by the C-CRT and the two-stage CRT with complex-valued moduli, or the closed-form CRTs with real-valued moduli. Note that to have an integer lcm, the complex-valued moduli can be selected as pairs of conjugate Gaussian integers.
@@
$\Delta_{\max}$ Method $M\mathsf{\Gamma}_{i}$ Dynamic range $M\Gamma$
@@ 9 C-CRT, Two-stage CRT $7+4\mathrm{i}$, $7-4\mathrm{i}$, 9 585
where $A$ is a constant, coefficients $a_{k}$ and $b_{k}$ are uniformly distributed in [1,1][-1,1], $\mathrm{sinc}(t)=\frac{\sin(\pi t)}{\pi t}$. For the closed-form CRTs, we set six SR-ADCs with dynamic ranges of $5,5,6,6,7$, and $7$ when $\Delta_{\max}=7$, and $7,7,8,8,9$, and $9$ when $\Delta_{\max}=9$. For the C-CRT and the two-stage CRT, we set six SR-ADCs with dynamic ranges of $\sqrt{41}$, $\sqrt{41}$, $\sqrt{41}$, $\sqrt{41}$, $7$ and $7$ when $\Delta_{\max}=7$, and $\sqrt{65}$, $\sqrt{65}$, $\sqrt{65}$, $\sqrt{65}$, $9$ and $9$ when $\Delta_{\max}=9$, where the moduli $M\mathsf{\Gamma}_{i}$ are shown in Table I. The reconstruction error is quantified using the root relative squared error (RRSE), given by
@@
where the sampling frequency $f_{s}$ in the time domain for each channel is the Nyquist rate, i.e., 11Hz.
Fig. 13 illustrates the RRSE curves for the three methods in terms of SNR. The fast MLE C-CRT demonstrates the best performance overall. Similar to the previous RMSE simulations, the fast MLE C-CRT achieves robustness at an SNR of 1616dB. In contrast, the two-stage CRT achieves robustness at an SNR of 1818dB. Additionally, for SNR values of 1818dB and higher, the performances of the two-stage CRT and the closed-form CRT are nearly indistinguishable.
@@
To show the high dynamic range of complex moduli, while the maximum dynamic range of all the SR-ADCs is constrained by $\Delta_{\max}$, we compute the TFR of the robust reconstruction for the three methods. The curves of the TFR for the three methods in terms of SNR are illustrated in Fig. 14, where $a_{k}=-0.9$, $b_{k}=0.98$, and $\tau=0.25$. Since some sampled values exceed the dynamic range of SR-ADCs with real-valued moduli, the closed-form CRTs exhibit an error floor. Additionally, although the TFRs of both the fast MLE C-CRT and the two-stage CRT can approach 0, the fast MLE C-CRT outperforms the two-stage CRT due to its optimal estimation of the common remainder. These results illustrate that the proposed MLE C-CRT for complex numbers performs better than the conventional CRT for real values. In addition, if C-CRT is thought of as a special 2D-CRT, it clearly shows that 2D-CRT (non-separable) performs better than two 1D-CRTs (separable).